Q3. Solve for vo(t) in the circuit of Fig. 3 using the Superposition 62 2H 12 cos 31 V + 4 sin 21 A 10 V Solution:
Answer #1

62 27 000 W 120083t v 4sin(24) 12 10V Take Toy Sourla deactive other sources un 2.44 oon le 10 Here W=0 Under Do Xc: Jwe ->O-then Xe= JWC Jx3x T24 4-5-6) Xc and XL Ejwls 6 jx3x2 = 16 J6 12L09 Voe- v 62 16 12/09 Vog Voz x 1220° from noiteye devison suNoz 24x12/09 24 tj 12 2X12 Loo 24 j 510.733-2056° Voz 10.7334-260568 => Vo2 = 10.133 608 (37-26156) face row lastly take 4 sNo3= -48 Loº 2-j = 21.466/26.56° Vort) = 21.466 608/ 27 +26.56) Voz It) = 21.466 Sin.(2++26.56°) than Total Vo (+) voltage V
Answer #2

A 2000 kW turbine –generator of 0.85 power factor operates at the rated load. An

additional load of 300kW at 0.8 power factor is added. Determine the required kVAR

of capacitor to operate the turbine generator but keep it from being overloaded.


answered by: jay
Answer #3

image.png

answered by: jay